solve the following systems of equations and state if the lines are parallel:
y+4x+5=2
y+3x=2

Found 2 solutions by richwmiller, Alan3354:
Answer by richwmiller(9135)

subtract 5 from both sides in the first
y+4x=-3
y+3x=2
subtract second from first
x=-5

Answer by Alan3354(30993)

y+4x+5=2
y+3x=2
----------
y+4x=-3
y+3x=2
------- Subtract
x = -5
-------
y = 17
----------
They are not parallel, the slopes are not equal.